OpenStudy (anonymous):

Calculate d^2y/dx^2 y = - 4/x^2

OpenStudy (anonymous):

It's asking for the second derivative. y = -4 (x)^-2 y' = 8 x^-3 y'' = -24 x^-4 Your answer should be -24/(x^-4)

OpenStudy (anonymous):

i see, however i plugged your answer into my homework assignment and it said it was incorrect.

OpenStudy (anonymous):

The original problem was the second derivative of: \[y = -4/x ^{2}\] ?

OpenStudy (anonymous):

or was it \[(-4/x)^{2}\]

OpenStudy (anonymous):

y= -4/x^2

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

The answer is \[-24/x ^{4}\], I'm positive and so is Wolfram Alpha. Make sure you have any negative signs or exponents in the right place so that it's formatted correctly, that has to be the issue. Try -24/(x^4)

OpenStudy (anonymous):

Oh wow, my bad. Forgot to get switch to a positive exponent when I flipped the x variable to the denominator. The right answer is the most recent post.

OpenStudy (anonymous):

oh ok, i got it now, thanks